A random sample of n=70 measurements is obtained from a population with a mean 200 and standard deviation of 50.describe the sampling for the sample means by computing

1
Expert's answer
2021-05-18T12:24:05-0400

The Central Limit Theorem

For samples of size 30 or more, the sample mean is approximately normally distributed, with mean "\mu_{\bar{X}}=\mu" and standard deviation "\sigma_{\bar{X}}=\sigma/\sqrt{n}," where "n"  is the sample size.

The larger the sample size, the better the approximation.


"\bar{X}\sim N(\mu, \sigma^2/n)"

Given "\mu=200, \sigma=50, n=70."


The distribution of the sample mean is normal with mean "\mu_{\bar{X}}=\mu=200" and standard deviation "\sigma_{\bar{X}}=\sigma/\sqrt{n}=50/\sqrt{70}"


"\bar{X}\sim N(200, 50^2/70)"